### v1

@@ Appendix B Regularity estimates in flat domains
Since $u$ is compactly supported on $\Omega$, we deduce from [35, Lemmas 9.16 and 9.17] that $u\in\mathrm{W}^{2,q}(\Omega)$ and
$\|u\|_{2,q;\Omega}\leq C\|Lu\|_{0,q;\Omega}=C\|f\|_{0,q;\Omega};$
this proves the case $m=0$. We now proceed by induction. Assume the Lemma holds true for a fixed $m\geq 0$. Consider now the setting given by $m$ replaced by $m+1$. Then, by the induction hypothesis $u\in\mathrm{W}^{m+2,q}(\Omega)$ and
$\|u\|_{m+2,q;\Omega}\lesssim\|f\|_{m,q;\Omega}.$
For each $i=1,\dots,d$, we differentiate (B.1) to deduce that $w_{i}:=\partial_{i}u\in\mathrm{W}^{m+2,p}(\Omega)$ (which is compactly supported on $\Omega$) satisfies
$\mathbf{A}:\nabla^{2}w_{i}+\mathbf{b}\cdot\nabla w_{i}+c\,w_{i}=g_{i},\qquad\text{a.e.~on $\Omega$},$
where $g_{i}:=\partial_{i}f-(\partial_{i}\mathbf{A}):\nabla^{2}u+(\partial_{i}\mathbf{b})\cdot\nabla u+(\partial_{i}c)\,u\in\mathrm{W}^{m,q}(\Omega)$. Then, by the induction hypothesis $w_{i}\in\mathrm{W}^{m+2,q}(\Omega)$ and
$\|\partial_{i}u\|_{m+2,q;\Omega}=\|w_{i}\|_{m+2,q;\Omega}\lesssim\|g_{i}\|_{m,q;\Omega}\lesssim\left(\|f\|_{m+1,q;\Omega}+\|u\|_{m+2,q;\Omega}\right)\lesssim\|f\|_{m+1,q;\Omega}.$
Consequently,
$\|u\|_{m+3,q;\Omega}\lesssim\|u\|_{0,q;\Omega}+\sum_{i=1}^{d}\|\partial_{i}u\|_{m+2,q;\Omega}\lesssim\|f\|_{m+1,q;\Omega},$
This finishes the induction step and, hence, the proof. ∎
Appendix C Calculus identities on manifolds
In this section we present some results on tangential calculus. Our main objective is to explicitly establish the relation between the Laplace–Beltrami $\Delta_{\Gamma}$ and Bochner Laplacian $\Delta_{B}$ operators, which proves essential in Section 3.
Within this section we adopt the Einstein summation convention.
For $\mathbf{u}\in\mathbf{C}^{2}_{t}(\Gamma)$ we define $\Delta_{\Gamma}\mathbf{u}$ to be the vector field given by the componentwise application of the Laplace–Beltrami operator on $\mathbf{u}$, i.e. $(\Delta_{\Gamma}\mathbf{u})_{i}=\Delta_{\Gamma}v_{i}$ for each $i=1,\dotsc,d+1$.
Proposition C.1 (normal part of the componentwise Laplace–Beltrami operator).
Suppose $\Gamma$ is of class $C^{3}$. Let $\mathbf{N}:=\mathbf{I}-\mathbf{P}=\boldsymbol{\nu}\otimes\boldsymbol{\nu}$. Then,
$\mathbf{N}\Delta_{\Gamma}\mathbf{u}=-(2\mathbf{B}:\nabla_{M}\mathbf{u}+\operatorname{div}_{\Gamma}\mathbf{B}\cdot\mathbf{u})\boldsymbol{\nu},\qquad\forall\mathbf{u}\in\mathbf{C}^{2}_{t}(\Gamma).$ (C.1)
@@ Appendix C Calculus identities on manifolds
We start noticing that identity (3.25) can be equivalently stated componentwise as $B_{qr}u_{r}=-\underline{D}_{q}u_{r}\nu_{r}$. Differentiating both sides of the identity leads to
$(\underline{D}_{s}\underline{D}_{q}u_{r})\nu_{r}=-\underline{D}_{q}u_{r}\underline{D}_{s}\nu_{r}-\underline{D}_{s}B_{qr}u_{r}-B_{qr}\underline{D}_{s}u_{r}=-\underline{D}_{q}u_{r}B_{sr}-\underline{D}_{s}B_{qr}u_{r}-B_{qr}\underline{D}_{s}u_{r},$
Hence,
$\Delta_{\Gamma}\mathbf{u}\cdot\boldsymbol{\nu}=(\underline{D}_{k}\underline{D}_{k}u_{r})\nu_{r}=-\underline{D}_{k}u_{r}B_{kr}-\underline{D}_{k}B_{kr}u_{r}-B_{kr}\underline{D}_{k}u_{r}=-2\mathbf{B}:\nabla_{M}\mathbf{u}-\operatorname{div}_{\Gamma}\mathbf{B}\cdot\mathbf{u},$
which completes the proof. ∎
Proposition C.2 (relation between the Bochner Laplacian and the componentwise Laplace–Beltrami operator).
Suppose $\Gamma$ is of class $C^{3}$. Then
$\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{u}=\Delta_{\Gamma}\mathbf{u}+\mathbf{B}^{2}\mathbf{u}+(2\mathbf{B}:\nabla_{M}\mathbf{u}+\operatorname{div}_{\Gamma}\mathbf{B}\cdot\mathbf{u})\boldsymbol{\nu},\qquad\forall\mathbf{u}\in\mathbf{C}^{2}_{t}(\Gamma).$ (C.2)
Proof.
Recalling definitions (2.1), (2.2) and (2.6), the fact that $\mathbf{P}^{2}=\mathbf{P}$, we initially obtain
$(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{u})_{i}$ $=P_{ij}(\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{u})_{j}=P_{ij}\underline{D}_{k}(\nabla_{\Gamma}\mathbf{u})_{jk}$
$=P_{ij}\underline{D}_{k}(\mathbf{P}\nabla_{M}\mathbf{u})_{jk}=P_{ij}\underline{D}_{k}(P_{jl}(\nabla_{M}\mathbf{u})_{lk})$
$=P_{ij}\underline{D}_{k}(P_{jl}\underline{D}_{k}u_{l})=P_{ij}\underline{D}_{k}P_{jl}\underline{D}_{k}u_{l}+P_{ij}P_{jl}\underline{D}_{k}\underline{D}_{k}u_{l}.$
Moreover recall that $\mathbf{B}\mathbf{P}=\mathbf{B}$, $\mathbf{P}\boldsymbol{\nu}=\mathbf{0}$, and since $\mathbf{P}_{jl}=\delta_{jl}-\nu_{j}\nu_{l}$ and $\mathbf{B}=\nabla_{M}\boldsymbol{\nu}$, it follows that $\underline{D}_{k}P_{jl}=-B_{kj}\nu_{l}-B_{kl}\nu_{j}$; hence,
$(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{u})_{i}$ $=-P_{ij}B_{kj}\nu_{l}\underline{D}_{k}u_{l}-P_{ij}B_{kl}\nu_{j}\underline{D}_{k}u_{l}+P_{il}\underline{D}_{k}\underline{D}_{k}u_{l}$ (C.3)
$=P_{il}\underline{D}_{k}\underline{D}_{k}u_{l}-B_{ik}\nu_{l}\underline{D}_{k}u_{l}=(\mathbf{P}\Delta_{\Gamma}\mathbf{u})_{i}-(\mathbf{B}\nabla_{M}^{T}\mathbf{u}\,\boldsymbol{\nu})_{i}=(\mathbf{P}\Delta_{\Gamma}\mathbf{u})_{i}+(\mathbf{B}^{2}\mathbf{u})_{i},$
where in the last equality we have used identity (3.25). Combining identities (C.1) and (C.3) yields (C.2) and finishes the proof. ∎
Proposition C.3 (weak relation between the Bochner Laplacian and the componentwise Laplace–Beltrami operator).
Let $1<p<\infty$ and let $p^{*}:=\frac{p}{p-1}$ be the dual exponent of $p$. Suppose $\Gamma$ is of class $C^{3}$. Then,
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}(\mathbf{P}\mathbf{v})$ $=\int_{\Gamma}\nabla_{\Gamma}u_{i}\cdot\nabla_{\Gamma}v_{i}-\int_{\Gamma}\mathbf{B}\mathbf{u}\cdot\mathbf{B}\mathbf{v}-\int_{\Gamma}(2\mathbf{B}:\nabla_{M}\mathbf{u}+\operatorname{div}_{\Gamma}\mathbf{B}\cdot\mathbf{u})(\mathbf{v}\cdot\boldsymbol{\nu}),$ (C.4)
$=\int_{\Gamma}\nabla_{\Gamma}u_{i}\cdot\nabla_{\Gamma}v_{i}+2\int_{\Gamma}\mathbf{u}\cdot\operatorname{div}_{\Gamma}((\mathbf{v}\cdot\boldsymbol{\nu})\mathbf{B})-\int_{\Gamma}\mathbf{B}\mathbf{u}\cdot\mathbf{B}\mathbf{v}-\int_{\Gamma}(\mathbf{v}\cdot\boldsymbol{\nu})\operatorname{div}_{\Gamma}\mathbf{B}\cdot\mathbf{u},$
for all $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma),\mathbf{v}\in\mathbf{W}^{1,p^{*}}(\Gamma)$.
Proof.
Multiplying (C.2) by $\mathbf{v}\in\mathbf{C}^{1}(\Gamma)$ and utilizing integration-by-parts formulas [10, Lemma 1] and [11, Ap. B] (notice that $\mathbf{B}\boldsymbol{\nu}=\mathbf{0}$) yield (C.4) for $\mathbf{u}\in\mathbf{C}^{2}_{t}(\Gamma)$ and $\mathbf{v}\in\mathbf{C}^{1}(\Gamma)$. Then, the result follows by density (cf. 2.5 and 2.11). ∎
Proposition C.4.
Suppose $\Gamma$ is of class $C^{3}$. Then,
$\Delta_{\Gamma}(\mathbf{P}\mathbf{u})=\mathbf{P}\Delta_{\Gamma}\mathbf{u}-2(\mathbf{B}:\nabla_{M}\mathbf{u})\boldsymbol{\nu}-(\mathbf{u}\cdot\boldsymbol{\nu})\operatorname{div}_{\Gamma}\mathbf{B}-(\operatorname{div}_{\Gamma}\mathbf{B}\cdot\mathbf{u})\boldsymbol{\nu}-2\mathbf{B}\nabla^{T}_{M}(\mathbf{N}\mathbf{u})\,\boldsymbol{\nu},\qquad\forall\mathbf{u}\in\mathbf{C}^{2}(\Gamma).$ (C.5)
Consequently,
$\mathbf{P}\Delta_{\Gamma}(\mathbf{P}\mathbf{u})=\mathbf{P}\Delta_{\Gamma}\mathbf{u}-(\mathbf{u}\cdot\boldsymbol{\nu})\mathbf{P}\operatorname{div}_{\Gamma}\mathbf{B}-2\mathbf{B}\nabla^{T}_{M}(\mathbf{N}\mathbf{u})\,\boldsymbol{\nu},\qquad\forall\mathbf{u}\in\mathbf{C}^{2}(\Gamma).$ (C.6)
Proof.
Recalling the definition of $\Delta_{\Gamma}$, the facts that $B_{ij}=\underline{D}_{j}\nu_{i}=\underline{D}_{i}\nu_{j}$ and $\underline{D}_{k}P_{jl}=-B_{kj}\nu_{l}-B_{kl}\nu_{j}$, and using the product rule for covariant derivatives, we obtain after a series of elementary computations that
$(\Delta_{\Gamma}(\mathbf{P}\mathbf{u}))_{i}=\Delta_{\Gamma}(\mathbf{P}\mathbf{u})_{i}=\underline{D}_{k}\underline{D}_{k}(P_{ij}u_{j})=\underline{D}_{k}\left(\underline{D}_{k}P_{ij}u_{j}+P_{ij}\underline{D}_{k}u_{j}\right)=\underline{D}_{k}\left(-B_{ki}\nu_{j}u_{j}-B_{kj}\nu_{i}u_{j}+P_{ij}\underline{D}_{k}u_{j}\right)\\ =-\underline{D}_{k}B_{ki}\nu_{j}u_{j}-B_{ki}\underline{D}_{k}\nu_{j}u_{j}-B_{ki}\nu_{j}\underline{D}_{k}u_{j}-\underline{D}_{k}B_{kj}\nu_{i}u_{j}-B_{kj}\underline{D}_{k}\nu_{i}u_{j}-B_{kj}\nu_{i}\underline{D}_{k}u_{j}+\underline{D}_{k}P_{ij}\underline{D}_{k}u_{j}+P_{ij}\underline{D}_{k}\underline{D}_{k}u_{j}\\ =-\underline{D}_{k}B_{ki}\nu_{j}u_{j}-B_{ki}B_{kj}u_{j}-B_{ki}\nu_{j}\underline{D}_{k}u_{j}-\underline{D}_{k}B_{kj}\nu_{i}u_{j}-B_{kj}B_{ki}u_{j}-B_{kj}\nu_{i}\underline{D}_{k}u_{j}-B_{ki}\nu_{j}\underline{D}_{k}u_{j}-B_{kj}\nu_{i}\underline{D}_{k}u_{j}+P_{ij}\underline{D}_{k}\underline{D}_{k}u_{j}.$
In turn, the definitions of the covariant divergence $\operatorname{div}_{\Gamma}$ and extrinsic gradient $\nabla_{M}$ (cf. Section 2.1), let us express the right-hand side in a more compact way, thereby obtaining
@@
$(\Delta_{\Gamma}(\mathbf{P}\mathbf{u}))_{i}=\left(\mathbf{P}\Delta_{\Gamma}\mathbf{u}-2(\mathbf{B}:\nabla_{M}\mathbf{u})\boldsymbol{\nu}-(\mathbf{u}\cdot\boldsymbol{\nu})\operatorname{div}_{\Gamma}\mathbf{B}-(\operatorname{div}_{\Gamma}\mathbf{B}\cdot\mathbf{u})\boldsymbol{\nu}-2\mathbf{B}^{2}(\mathbf{P}\mathbf{u})-2\mathbf{B}\nabla^{T}_{M}\mathbf{u}\,\boldsymbol{\nu}\right)_{i}.$
In turn, we can rewrite the second to last term in the right-hand side with the help of identity (3.25) applied to $\mathbf{P}\mathbf{u}\in\mathbf{C}^{2}_{t}(\Gamma)$, thereby arriving at
@@
$(\Delta_{\Gamma}(\mathbf{P}\mathbf{u}))_{i}$ $=\left(\mathbf{P}\Delta_{\Gamma}\mathbf{u}-2(\mathbf{B}:\nabla_{M}\mathbf{u})\boldsymbol{\nu}-(\mathbf{u}\cdot\boldsymbol{\nu})\operatorname{div}_{\Gamma}\mathbf{B}-(\operatorname{div}_{\Gamma}\mathbf{B}\cdot\mathbf{u})\boldsymbol{\nu}+2\mathbf{B}\nabla^{T}_{M}(\mathbf{P}\mathbf{u})\,\boldsymbol{\nu}-2\mathbf{B}\nabla^{T}_{M}\mathbf{u}\,\boldsymbol{\nu}\right)_{i}$

### v2

@@ Appendix B Calculus identities on manifolds
Let $\mathsf{d}$ denote the signed distance function [10, Section 1.2.3] defined on a sufficiently small tubular neighborhood of $\Gamma$, which we assume without loss of generality to coincide with $\Omega_{\delta}$. Since $\Gamma$ is of class $C^{2}$, we know that $\mathsf{d}\in C^{2}(\Omega_{\delta})$. The closest point projection $\mathbf{P}_{\mathsf{d}}:\Omega_{\delta}\rightarrow\mathrm{R}$ is defined by $\mathbf{P}_{\mathsf{d}}(\mathbf{x})=\mathbf{x}-\mathsf{d}(x)\nabla\mathsf{d}(\mathbf{x})$ for each $\mathbf{x}\in\Omega_{\delta}$. Notice that $\nabla\mathbf{P}_{\mathsf{d}}=\mathbf{P}\circ\mathbf{P}_{\mathsf{d}}-\mathsf{d}\nabla^{2}\mathsf{d}$ in $\Omega_{\delta}$. Let the superindex n denote the constant normal extension [10, eq. (1.26)] of functions defined on $\Gamma$ to $\Omega_{\delta}$. More precisely, for each $v:\Gamma\rightarrow\mathrm{R}$, the function $v^{n}:\Omega_{\delta}\rightarrow\mathrm{R}$ is defined by $v^{n}:=v\circ\mathbf{P}_{\mathsf{d}}$. Since $\Gamma$ is of class $C^{2}$, then for each $v\in C^{2}(\Gamma)$ we can only ensure $v^{n}:\Omega_{\delta}\rightarrow\mathrm{R}$ to be of class $C^{1}$. This is because differentiating $v^{n}=v\circ\mathbf{P}_{\mathsf{d}}$ amounts to differentiating $\mathbf{P}_{\mathsf{d}}=\mathrm{id}-\mathsf{d}\nabla\mathsf{d}$, which already contains $\nabla\mathsf{d}$. From [10, p. 15] we know that the extensions e and n are related with the differential operators $\nabla_{M}$ on $\Gamma$ acting on scalar and vector-valued functions $v$ and $\mathbf{v}$ via the identity
$\nabla_{M}v=(\nabla v^{n})\rvert_{\Gamma}=\mathbf{P}(\nabla v^{e})\rvert_{\Gamma},\qquad\nabla_{M}\mathbf{v}=(\nabla\mathbf{v}^{n})\rvert_{\Gamma}=(\nabla v^{e})\rvert_{\Gamma}\mathbf{P},\qquad\text{on $\Gamma$}.$
Recall also that $\nabla_{\Gamma}v=\nabla_{M}v$ and $\nabla_{\Gamma}\mathbf{v}=\mathbf{P}\nabla_{M}\mathbf{v}$.
By the chain rule, we have that
$\nabla[(\nabla v^{e})\rvert_{\Gamma}]^{n}=\nabla\left((\nabla v^{e})\circ\mathbf{P}_{\mathsf{d}}\right)=\left((\nabla^{2}v^{e})\circ\mathbf{P}_{\mathsf{d}}\right)\nabla\mathbf{P}_{\mathsf{d}},\qquad\text{in $\Omega_{\delta}$},$
which restricted to $\Gamma$ reduces to
$\nabla_{M}\left((\nabla v^{e})\rvert_{\Gamma}\right)=(\nabla^{2}v^{e})\rvert_{\Gamma}\mathbf{P},\qquad\text{on $\Gamma$}.$
Moreover, by multiplying by $\mathbf{P}$ from the left, we also obtain
[…]
Recalling that $\mathbf{P}+\mathbf{N}=\mathbf{I}$, we can decompose $(\nabla v^{e})\rvert_{\Gamma}$ as $(\nabla v^{e})\rvert_{\Gamma}=\mathbf{P}(\nabla v^{e})\rvert_{\Gamma}+\mathbf{N}(\nabla v^{e})\rvert_{\Gamma}=\nabla_{\Gamma}v+\mathbf{N}(\nabla v^{e})\rvert_{\Gamma}$. Moreover, using the first identity of B.1 with $\mathbf{v}=(\nabla v^{e})\rvert_{\Gamma}$ we have that $\nabla_{\Gamma}(\mathbf{N}(\nabla v^{e})\rvert_{\Gamma})=\nabla_{\Gamma}((\nabla v^{e})\rvert_{\Gamma})-\nabla_{\Gamma}(\mathbf{P}(\nabla v^{e})\rvert_{\Gamma})=((\nabla v^{e})\rvert_{\Gamma}\cdot\boldsymbol{\nu})\mathbf{B}$. Consequently, (B.1) becomes
$\nabla_{\Gamma}\nabla_{\Gamma}v+((\nabla v^{e})\rvert_{\Gamma}\cdot\boldsymbol{\nu})\mathbf{B}=\mathbf{P}(\nabla^{2}v^{e})\rvert_{\Gamma}\mathbf{P},$
which is precisely (2.2). ∎
The symmetry of the covariant Hessian allows us to provide a simple proof of the commutator identity (2.3).
Proposition B.3.
(Dziuk–Elliott’s commutator identity) Suppose $\Gamma$ is of class $C^{2}$. Then, for each $v\in C^{2}(\Gamma)$ and for each $i,j\in\{1,\dotsc,d+1\}$, identity (2.3) holds true. That is,
@@
First, notice that identity (2.3) is equivalent to
$\nabla_{M}^{T}\nabla_{M}v-\nabla_{M}\nabla_{M}v=\boldsymbol{\nu}(\mathbf{B}\nabla_{M}v)^{T}-(\mathbf{B}\nabla_{M}v)\boldsymbol{\nu}^{T},$
which we now prove. Recalling identity $\mathbf{P}+\mathbf{N}=\mathbf{I}$, and relation $\nabla_{\Gamma}\mathbf{v}=\mathbf{P}\nabla_{M}\mathbf{v}$ for vector-valued functions $\mathbf{v}$, we rewrite the second term on the left-hand side as follows
$\nabla_{M}\nabla_{M}v=\nabla_{\Gamma}\nabla_{M}v+\mathbf{N}\nabla_{M}(\nabla_{M}v)=\nabla_{\Gamma}\nabla_{M}v-\boldsymbol{\nu}(\nabla_{M}^{T}v)\mathbf{B},$
where in the last equality we have used (2.7) with $\mathbf{v}=\nabla_{M}v\in\mathbf{C}^{1}_{t}(\Gamma)$ to write $\mathbf{N}\nabla_{M}(\nabla_{M}v)=-\boldsymbol{\nu}(\nabla_{M}^{T}v)\mathbf{B}$. Consequently, recalling from B.2 that $\nabla_{\Gamma}\nabla_{\Gamma}v=\nabla_{\Gamma}\nabla_{M}v$ is symmetric, we arrive at
$\nabla_{M}^{T}\nabla_{M}v-\nabla_{M}\nabla_{M}v=-\mathbf{B}(\nabla_{M}v)\boldsymbol{\nu}^{T}+\boldsymbol{\nu}(\nabla_{\Gamma}^{T}v)\mathbf{B}.$
This finishes the proof. ∎
We now prove the important 2.13 (weak commutator identity), that extends [36, Lemma 2.1].
Proof of 2.13.
Let us first assume that $\mathbf{v}\in\mathbf{C}^{2}(\Gamma)$ (not necessarily tangential) and $\phi\in C^{2}(\Gamma)$. Then, the symmetry of the covariant Hessian $\nabla_{\Gamma}\nabla_{\Gamma}\phi$ (2.2), integration-by-parts formula [11, Lemma B.7] and the fact that $\nabla_{\Gamma}^{T}\mathbf{v}\,\boldsymbol{\nu}=\textbf{0}$ yield
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{v}:\nabla_{\Gamma}\nabla_{\Gamma}\phi=-\int_{\Gamma}\nabla_{\Gamma}^{T}\mathbf{v}:\nabla_{\Gamma}\nabla_{\Gamma}\phi=-\int_{\Gamma}\nabla_{\Gamma}^{T}\mathbf{v}:\nabla_{M}\nabla_{\Gamma}\phi=\int_{\Gamma}\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}^{T}\mathbf{v}\cdot\nabla_{\Gamma}\phi.$
Let us carefully manipulate the quantities appearing in the right-hand side of this last equation. By means of the product rule and the commutator identity (2.3), we first write
$(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}^{T}\mathbf{v})_{i}$ $=P_{ik}(\operatorname{div}_{\Gamma}\nabla_{\Gamma}^{T}\mathbf{v})_{k}=P_{ik}\underline{D}_{l}(\nabla_{\Gamma}^{T}\mathbf{v})_{kl}$
$=P_{ik}\underline{D}_{l}(\nabla_{M}^{T}\mathbf{v}\,\mathbf{P})_{kl}=P_{ik}\underline{D}_{l}(\underline{D}_{k}v_{m}P_{ml})=P_{ik}(\underline{D}_{l}\underline{D}_{k}v_{m})P_{ml}+P_{ik}(\underline{D}_{k}v_{m})(\underline{D}_{l}P_{ml})$
[…]
which by recalling that $\mathbf{P}=\mathbf{I}-\boldsymbol{\nu}\otimes\boldsymbol{\nu}$ and $\mathbf{B}\boldsymbol{\nu}=\textbf{0}$, and identities $\underline{D}_{k}P_{jl}=-B_{kj}\nu_{l}-B_{kl}\nu_{j}$ and $B_{qr}u_{r}=-\underline{D}_{q}u_{r}\nu_{r}$ (see the proof of B.1), can be further rewritten as follows:
$(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}^{T}\mathbf{v})_{i}$ $=(\underline{D}_{i}\underline{D}_{l}v_{m})P_{ml}-P_{ik}(\underline{D}_{k}v_{m})B_{ll}\nu_{m}-P_{ik}(\underline{D}_{k}v_{m})B_{ml}\nu_{l}$
$=(\underline{D}_{i}\underline{D}_{l}v_{m})P_{ml}-\operatorname*{tr}(\mathbf{B})(\underline{D}_{i}v_{m})\nu_{m}$
$=\underline{D}_{i}\underline{D}_{l}v_{l}-(\underline{D}_{i}\underline{D}_{l}v_{m})\nu_{m}\nu_{l}-\operatorname*{tr}(\mathbf{B})(\underline{D}_{i}v_{m})\nu_{m}.$
Whence, together with integration-by-parts formula [11, Lemma B.5], we get
@@
$\int_{\Gamma}\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}^{T}\mathbf{v}\cdot\nabla_{\Gamma}\phi=\int_{\Gamma}\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}\cdot\nabla_{\Gamma}\phi-\int_{\Gamma}(\underline{D}_{i}\underline{D}_{l}v_{m})\nu_{m}\nu_{l}\,\underline{D}_{i}\phi-\int_{\Gamma}\operatorname*{tr}(\mathbf{B})\left(\nabla_{M}^{T}\mathbf{v}\,\boldsymbol{\nu}\right)\cdot\nabla_{\Gamma}\phi\\ =-\int_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}\,\Delta_{\Gamma}\phi+\int_{\Gamma}(\underline{D}_{l}v_{m})\underline{D}_{i}(\nu_{m}\nu_{l}\,\underline{D}_{i}\phi)-\int_{\Gamma}\operatorname*{tr}(\mathbf{B})(\underline{D}_{l}v_{m})\nu_{m}\nu_{l}\cancel{(\underline{D}_{i}\phi)\nu_{i}}-\int_{\Gamma}\operatorname*{tr}(\mathbf{B})\left(\nabla_{M}^{T}\mathbf{v}\,\boldsymbol{\nu}\right)\cdot\nabla_{\Gamma}\phi\\ =-\int_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}\,\Delta_{\Gamma}\phi+\int_{\Gamma}(\underline{D}_{l}v_{m})(\underline{D}_{i}\nu_{l})\nu_{m}\,\underline{D}_{i}\phi+\int_{\Gamma}\cancel{(\underline{D}_{l}v_{m})\nu_{l}}\underline{D}_{i}(\nu_{m}\,\underline{D}_{i}\phi)-\int_{\Gamma}\operatorname*{tr}(\mathbf{B})\left(\nabla_{M}^{T}\mathbf{v}\,\boldsymbol{\nu}\right)\cdot\nabla_{\Gamma}\phi\\ =-\int_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}\,\Delta_{\Gamma}\phi+\int_{\Gamma}(\mathbf{B}\nabla_{M}^{T}v\,\boldsymbol{\nu})\cdot\nabla_{\Gamma}\phi-\int_{\Gamma}\operatorname*{tr}(\mathbf{B})\left(\nabla_{M}^{T}\mathbf{v}\,\boldsymbol{\nu}\right)\cdot\nabla_{\Gamma}\phi.$
So far, we have proved that for each $\phi\in C^{2}(\Gamma)$ and $\mathbf{v}\in\mathbf{C}^{2}(\Gamma)$, we have that
@@
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{v}:\nabla_{\Gamma}\nabla_{\Gamma}\phi=-\int_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}\,\Delta_{\Gamma}\phi+\int_{\Gamma}(\mathbf{B}\nabla_{M}^{T}\mathbf{v}\,\boldsymbol{\nu})\cdot\nabla_{\Gamma}\phi-\int_{\Gamma}\operatorname*{tr}(\mathbf{B})\left(\nabla_{M}^{T}\mathbf{v}\,\boldsymbol{\nu}\right)\cdot\nabla_{\Gamma}\phi;$
which, by means of a density argument, can be easily extended to $\mathbf{v}\in\mathbf{C}^{1}(\Gamma)$. In particular, by choosing $\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma)$ and using that $\mathbf{B}\mathbf{v}=-(\nabla_{M}^{T}\mathbf{v})\boldsymbol{\nu}$ (cf. (2.6)), we arrive at (2.8). This finishes the proof. $\square$
Appendix C Riemann curvature tensor
This section is devoted to extending the notion of the classical Riemann Curvature tensor [20] to a “weaker” setting. This is instrumental in the proof of 3.3 (Poincaré inequality in $\mathbf{W}^{1,p}_{t}(\Gamma)$).
Within this section we write $\partial_{\mathbf{a}}\mathbf{v}:=(\nabla_{M}\mathbf{v})\mathbf{a}$, $D_{\mathbf{a}}\mathbf{v}:=(\nabla_{\Gamma}\mathbf{v})\mathbf{a}$ and $\mathbf{a},\mathbf{b}]:=\partial_{\mathbf{a}}\mathbf{b}-\partial_{\mathbf{b}}\mathbf{a}$.
For $\Gamma$ of class $C^{3}$, the Riemann Curvature tensor [20, §4.2, Definition 2.1] $R$ is defined by
$R(\mathbf{a},\mathbf{b},\mathbf{v}):=D_{\mathbf{a}}D_{\mathbf{b}}\mathbf{v}-D_{\mathbf{b}}D_{\mathbf{a}}\mathbf{v}-D_{[\mathbf{a},\mathbf{b}]}\mathbf{v}\in\mathbf{C}_{t}(\Gamma),$
[…]
$R(\mathbf{a},\mathbf{b},\mathbf{v})=\mathbf{P}(\partial_{\mathbf{a}}\partial_{\mathbf{b}}\mathbf{v}-\partial_{\mathbf{b}}\partial_{\mathbf{a}}\mathbf{v}-\partial_{[\mathbf{a},\mathbf{b}]}\mathbf{v})+(\mathbf{v}^{T}\mathbf{B}\mathbf{b})\mathbf{B}\mathbf{a}-(\mathbf{v}^{T}\mathbf{B}\mathbf{a})\mathbf{B}\mathbf{b}.$ (C.1)
Furthermore, by taking into account the easily-verifiable identity
$(\partial_{\mathbf{a}}\partial_{\mathbf{b}}\mathbf{v})_{i}=\mathbf{a}^{T}\left\{(\nabla_{M}\nabla_{M}v_{i})\mathbf{b}+(\nabla_{M}^{T}\mathbf{b})(\nabla_{M}v_{i})\right\},$
and its analogue for $(\partial_{\mathbf{b}}\partial_{\mathbf{a}}\mathbf{v})_{i}$, and noticing that $\mathbf{a}^{T}(\nabla_{M}\nabla_{M}v_{i})\mathbf{b}=\mathbf{b}^{T}(\nabla_{M}\nabla_{M}v_{i})\mathbf{a}$ because of (2.3) and the tangentiality of $\mathbf{a}$ and $\mathbf{b}$, it is easy to see that $\partial_{\mathbf{a}}\partial_{\mathbf{b}}\mathbf{v}-\partial_{\mathbf{b}}\partial_{\mathbf{a}}\mathbf{v}=\partial_{[\mathbf{a},\mathbf{b}]}\mathbf{v}$; hence (C.1) reduces to
$R(\mathbf{a},\mathbf{b},\mathbf{v})=(\mathbf{v}^{T}\mathbf{B}\mathbf{b})\mathbf{B}\mathbf{a}-(\mathbf{v}^{T}\mathbf{B}\mathbf{a})\mathbf{B}\mathbf{b},\qquad\forall\mathbf{a},\mathbf{b}\in\mathbf{C}^{1}_{t}(\Gamma),\mathbf{v}\in\mathbf{C}^{2}_{t}(\Gamma).$ (C.2)
That is, $R$ reduces to a zeroth order differential operator. Our objective is now to extend $R$ to less regular manifolds and vector fields.
Indeed, assume from now on that $\Gamma$ is only of class $C^{2}$, integration-by-parts formulas [11, Ap. B] induce the following definition: for each $\mathbf{a},\mathbf{b},\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma)$ we define the “weak” Riemann Curvature tensor $R(\mathbf{a},\mathbf{b},\mathbf{v})$ as a linear functional acting on vector fields $\mathbf{w}\in\mathbf{C}^{1}_{t}(\Gamma)$ defined by